A piston has an external pressure of 5.00 atm. How much work has been done if the cylinder goes from a volume of 0.140 liters to 0.570 liters.
Answer #1

Work = pdeltav

Volume change = 0.57-0.14= 0.43L

Pressure= 5 atm

Work = 5*0.43=2.15